Ranking Investment Proposals: Payback Period, Accounting Rate of Return, and Net Present Value
Presented is information pertaining to the cash flows of three mutually exclusive investment proposals:

Proposal X Proposal Y Proposal Z
Initial investment $98,000 $98,000 $98,000
Cash flow from operations
Year 1 90,000 49,000 98,000
Year 2 8,000 49,000
Year 3 49,000 49,000
Disinvestment 0 0 0
Life (years) 3 years 3 years 1 year

(a) Select the best investment proposal using the payback period, the accounting rate of return on initial investment, and the net present value criteria. Assume that the organization's cost of capital is 10 percent.

  • Round accounting rate of return four decimal places.

  • Round net present value to the nearest whole number.

  • Use negative signs with your answers, when appropriate.

Proposal X Proposal Y Proposal Z Best proposal
Payback period (years) Answer Answer Answer AnswerXYZX,YX,ZY,Z
Accounting rate of return Answer Answer Answer AnswerXYZX,YX,ZY,Z
Net present value Answer Answer Answer AnswerXYZX,YX,ZY,Z
Answer #1

Propasal X

Net present value=PV of cash inflow - PV of cash outflow

=90000*PVF(10% , n year)+8000*PVF(10%, n year)+49000*PVF(10%, n year) - 98000

=90000*909+8000*.826+49000*.751-98000

=$27217

Payback Period

Project X

Year

Cash Flow

Cumulative Cash Flow

0

(98000)

(98000)

1

90,000

(8000)

2

8,000

-

3

49,000

49,000

=1+(8000/8000)

= 2 years

Accounting rate of return = Averate Income / Average Investment

Average Income = (90000+8000+49000)/3 = 49000

ARR = 49000/98000

= 50%

Proposal Y

Net Present value = PV of cash inflow - PV of cash outflow

=49000*.909+49000*.826+49000*.751 - 98000

=$23814

Payback Period = Initial Investment / Cash inflow for period

98000 / 49000 = 2 years

Accounting rate of return = Average Income / Average Investment

Average Income= (49000+49000+49000)/3 = 49000

ARR = 49000 / 98000

= 50%

Proposal Z

Net present value = PV of cash inflow - PV of cash outflow

=98000*.909 = 98000

= - 8918

Payback period = Initial Investment / Cash inflow per period

98000/98000 = 1 year

Accounting rate of return = Average Income / Average Investment

= 98000/98000 = 100%

Proposal X should be accepted